OpenStudy (anonymous):

What is the discriminant of 3x2 + 6x = 2?

OpenStudy (jack1):

quadratic equation is ax^2 + bx +c = 0 ur equation is 3x^2 + 6x = 2.... so 3x^2 + 6x - 2 = 0 follow ok so far man?

OpenStudy (jack1):

@brosmean21 ... u there man?

OpenStudy (anonymous):

yeah good so far.

OpenStudy (jack1):

cool discriminant is always found using equation b^2 - 4ac

OpenStudy (jack1):

in ur equation: b=6 a = 3 c = (-2) so can u solve from here?

OpenStudy (anonymous):

added all up i got 60.

OpenStudy (jack1):

perfect

OpenStudy (anonymous):

thanks.

OpenStudy (jack1):

np ;)